# Mamba State-Space Models Are Lyapunov-Stable Learners

### v2

@@ 3.2 Hardware-aware optimizations and PEFT
Consider the weight matrix $\mathbf{W}$ of a MambaBlock from Equation 4. Targeting $\mathbf{W}$ for LoRA during fine-tuning ties adaptation weights across $\mathbf{B}_{t},\mathbf{C}_{t}$ and $\bm{\Delta}_{t}$.
The proof of Theorem 3 is available in Appendix A. The specific affects of both targeting $\mathbf{W}$ for fine-tuning and Theorem 3’s impact on generalization are ablated in Appendix F.
4 Related Work
Recent work has sought to understand how to efficiently increase Mamba’s hidden-state dimension by restructuring SSM operations using tensor contractions [11], leading to Mamba-2 models. A separate line of work has sought to understand the in-context learning (ICL) capabilities of Mamba LLMs when trained from scratch for specific tasks [40, 30]. Another line of recent work has sought to understand how hybrid Mamba-Transformer models may be directly distilled from Transformer models [54]. However, to the best of our knowledge, no existing works have either theoretically explored the effects small input changes (e.g., due to mixed-precision) have on Mamba’s recurrent dynamics, empirically explored such effects downstream impact on fine-tuning and inference, or sought to understand the effects of LoRA adaptation on modules within the MambaBlock.
Lyapunov exponents have previously been considered for classic RNN structures (e.g., vanilla RNNs, LSTMs, GRUs, PLRNNs, etc.) [39, 52], to determine when such models exhibit chaotic dynamics and the impact on the exploding/vanishing gradient phenomena***We note that this continues a long line of research exploring RNNs sensitivity to initial conditions and their subsequent ability to produce chaotic output [44, 29, 2, 3], although previous work did not leverage Lyapunov exponents.. For more recent S4 neural models, [18] used Hurwitz matrices to characterize the numerical stability of linear time-invariant (LTI) S4 models. However, such analysis is not applicable to time-varying models, such as Mamba, nor does it characterize the effects of sensitive dependence on initial conditions (e.g., divergence of two $\varepsilon$ close inputs). To the best of our knowledge, no previous works have used Lyapunov exponents to explore the effects of mixed-precision on recurrent neural models or Mamba architectures.
As previously noted, recent works [40, 30] have studied Mamba’s ability to perform ICL by training Mamba models for specific tasks. Such tasks include logistic regression, decision trees, and learning other simple function classes, following the work of [17]. We emphasize that, in this set up, relatively small Mamba models–33 million and 90 million parameters for [30] and [40], respectively–are trained from scratch for every evaluated task. Indeed, [40] notes that subsequent work is necessary to understand Mamba’s ICL capabilities for language modeling using standard natural language benchmarks, as well as for larger model sizes. Thus, our study of both the pretrained and instruction tuned ICL capabilities of Mamba/Mamba-2 LLMs for natural language tasks are complimentary to previous works.
@@ 5 Experiments
To demonstrate the implications of Theorem 2, we explore the performance difference between running inference with full-precision pretrained weights and using mixed-precision (FP16 and BF16) weights. Model performance is measured as percent accuracy using the MMLU dataset [24]. The difference in model performance is reported as the mean divergence (i.e., absolute difference) between the original full-precision and respective mixed-precision model, averaged over {0, 1, 3, 5}-shot percent accuracy. Thus, a divergence greater than one denotes an average difference greater than one entire percentage of accuracy.
Mamba pretrained checkpoints are compared to pretrained Transformer models of similar parameter counts and no more than $\sim$300B total pretraining tokens (Pythia [4], OLMo [19] 336B-token checkpoint, and Phi 1.5 [33]). We note that Pythia and Mamba models were both pretrained using the same corpus [15], allowing the fairest comparison between SSMs and Transformers. To limit extraneous numerical effects within experiments (e.g., due to parameter aggregation across multiple GPUs), all models were run using a single GPU (Nvidia A10G, 24 GB total memory). All models were evaluated using the LM evaluation harness from Eleuther AI [16]. Further experimental details are available in Appendix D. The results are available in Table 1.
Table 1: Mean full-precision (FP32) divergence in MMLU performance for mixed-precision inference. Divergence is averaged over {0, 1, 3, 5}-shot performance. Pretrained checkpoints are used for Mamba (M), Pythia (P), OLMo [19], and Phi-1.5 [33] (Phi) models.
Model M P M P M P OLMo M P Phi M P
Size 130M 160M 370M 410M  790M 1B 1.4B 1.5B 2.8B
FP16 $\mu$ 0.03 0.35 0.05 0.06 0.21 0.05 0.04 0.04 0.07 0.03 0.15 0.12
BF16 $\mu$ 0.05 1.45 0.20 0.20 0.66 0.16 0.13 0.31 0.13 1.05 1.17 0.11
From Table 1, inferencing in Pythia using FP16 and BF16 result in an average 0.13 and 0.41 full-precision divergence, respectively. Mamba displays similar averages in comparison: inferencing in Mamba using FP16 and BF16 result in an average 0.10 and 0.48 divergence, respectively. Interestingly, both SSM and Transformer architectures exhibit large divergence spikes–i.e., mean divergence greater than a percentage point–when using BF16, which occurs once for Mamba and Phi 1.5 models and twice for Pythia models. Due to space constraints, Mamba-2 results for the same experiment are included in Appendix Table 3. We note that, for comparable model sizes and mixed-precision, Mamba-2 models follow identical divergence trends as respective Mamba models.
In the following, we show that the observed large divergence spikes may be mitigated for Mamba SSMs by combining mixed-precision with parameter-efficient adapters during fine-tuning.
5.1 Mamba SSMs are stable learners
We next explore the implications of Theorem 1 on fine-tuning, wherein mixed-precision is especially critical; MPFT combined with PEFT adapters have been shown to drastically reduce Transformer fine-tuning times [12]. We are thus interested in the divergence between Mamba models fully fine-tuned (i.e., no adapters, all model weights are trained) in full-precision and models fine-tuned using mixed-precision and/or PEFT adapters. We focus on utilizing LoRA [25], which is arguably the most widely used PEFT framework for LLMs.
Figure 1: Mean full-precision (FP32) divergence in MMLU performance for Mamba, Pythia, and OpenELM models. Models are fine-tuned over the Alpaca dataset [48] using different combinations of MPFT and PEFT. Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
@@ 5.1 Mamba SSMs are stable learners
Using the Alpaca dataset [48], Mamba 160M, 410M, and 790M models are fine-tuned for three epochs with a maximum sequence length of 512. We denote the targeting of all linear layers (ALL) for LoRA as ALL LoRA, the targeting of a subset of linear layers (SLL) for LoRA as SLL LoRA, and no adapters as Full (i.e., full fine-tuning). Both ALL and SLL LoRA adapt the large memory buffer described in Theorem 3.
@@
Each fine-tuning run occurred on a single A10G GPU. To further limit extraneous numerical effects, the same batch size is used for all FP32, FP16, and BF16 experiments for a given model size. While this leads to hardware underutilization (i.e., non-saturated GPU memory for mixed-precision and LoRA experiments), this is necessary to guarantee no divergence is due to differences in parameter update schedules. For comparison, two Transformer-based LLM families of similar parameter counts are fine-tuned using the same experimental setup: Pythia (sizes 160M, 410M, and 1B) and OpenELM [37] (sizes 270M and 450M). The training recipe for all models was adapted from [50], with the AdamW_torch optimizer and a cosine annealing schedule. Further experimental details are available in Appendix D.
For each Mamba, Pythia, and OpenELM model, Figure 1 shows the mean divergence calculated between the respective FP32 Full and mixed-precision ALL/SLL LoRA fine-tuned models, averaged over {0, 1, 3, 5}-shot MMLU accuracy.
@@
Across mixed-precisions and adapter settings, Mamba displays smaller divergences than both Pythia and OpenELM models. E.g., for FP16, Mamba demonstrates an average divergence of 0.1, compared to 0.14 for Pythia and 0.54 for OpenELM. Similarly, for BF16, Mamba demonstrates an average divergence of 0.18, compared to 0.28 for Pythia and 0.33 for OpenELM. Importantly, Mamba models do not exhibit large deviation spikes after fine-tuning, in contrast to both Pythia and OpenELM models.
(a) Average tokens-per-second $\uparrow$.
(b) Maximum memory-per-token $\downarrow$.
Hardware throughput and memory-utilization improvements. With stable dynamics and observed divergences smaller than comparable Transformers, we show that MPFT and PEFT may be used to significantly increase GPU-training throughput for Mamba SSMs. To demonstrate such improvements, we utilize the previous fine-tuning settings for the Alpaca dataset. However, we now adjust the batch size to maximize throughput per MPFT and PEFT configuration.
For each MPFT and PEFT configuration, the average tokens-per-second (ATPS) is calculated as the total tokens used for fine-tuning divided by total training time, and the maximum memory-per-token (MMPT) is calculated as the maximum GPU memory utilization incurred (over the entire fine-tuning run) divided by the total number of tokens in each mini-batch. Results are plotted in Figure 2(b).
Both throughput and memory utilization improve as the number of Mamba parameters increases in Figure 2(b). Compared to the full-precision full fine-tuning of Mamba 790M (the largest model supported by an A10G’s memory capacity), evaluated MPFT and PEFT combinations result in an average 2.15 times more training tokens-per-second while reducing per-token memory utilization by an average 62.7%. Across all model sizes, evaluated MPFT and PEFT combinations result in an average 1.74 times more training tokens-per-second while reducing per-token memory utilization by an average 47.2% compared to respective full-precision fine-tuned runs. Furthermore, while full fine-tuning is no longer possible on a single A10G for Mamba models greater than 790 million parameters, MPFT and PEFT allow training Mamba models up to 2.8 billion parameters on GPUs with as little as 24 GB onboard memory.
@@
5.2 Instruction tuning impact on Mamba ICL for natural language tasks
Using both MPFT and PEFT, we next explore how instruction tuning affects Mamba and Mamba-2 ICL performance on natural language tasks. All Mamba and Mamba-2 pretrained models are instruction fine-tuned using ALL LoRA and the OpenHermes dataset [49] (which consists of 242,000 supervised samples). We use the training recipe of  [50], which includes BF16 utilization.
@@ 5.2 Instruction tuning impact on Mamba ICL for natural language tasks
Zero and few-shot performance is evaluated using five standard natural language benchmarks: HellaSwag [58], PIQA [5], Arc-E [8], Arc-C [8], and WinoGrande [46]. ICL performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot (AIPSS). For comparison, Pythia pretrained models are instruction fine-tuned using the same training recipe and ALL LoRA (i.e., all Pythia linear layers are adapted).
@@
Figure 3: Instruction tuning narrows the ICL gap between Mamba and Pythia, and creates a gap from Pythia to Mamba-2 models. ALL LoRA models were instruction tuned on the OpenHermes [49] dataset for one epoch. Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks.
@@
However, after instruction tuning, all Mamba models larger than Mamba 130M consistently improve in ICL performance as the number of shots increase. Similarly, the majority of Mamba-2 models larger than Mamba 130M greatly improve in ICL performance. Thus, while pretrained Mamba and Mamba-2 models are only capable of 38% and 82% (respectively) of the AIPSS compared to similar pretrained Pythia models, instruction tuned Mamba and Mamba-2 models are capable of 81.5% and 133% of the AIPSS relative to similarly fine-tuned Pythia models. We note that a significant difference between Mamba and Mamba-2 models is the larger (by a factor of four) latent dimension.
5.3 ICL is an emergent ability of instruction tuned Mamba SSMs
We next study the emergent behavior (as a function of model size) of Mamba/Mamba-2 SSMs’ ICL abilities on natural language tasks by comparing to a larger number of Transformer-based LLMs of varying sizes. We compare to OpenELM [37] (sizes 270M, 450M, and 1.1B), TinyLlama 1.1B [59], and OLMo 1.2B [19]. To limit the emergent effects on both parameter size and pretraining token counts, we did not evaluate models greater than 2.8 billion parameters and chose open-source checkpoints as close as possible to the 300 billion total pretraining tokens used for Mamba, Mamba-2, and Pythia models. Thus, pretraining token counts for OpenELM, TinyLlama, and OLMo models were 429 billion, 503 billion, and 336 billion, respectively. We note this potentially biases ICL performance in favor of the newly evaluated Transformer-based LLMs, and that direct comparisons between Mamba, Mamba-2, and Pythia are the most fair (as these three classes of models were all pretrained on the same dataset for the same number of total pretraining tokens).
[…]
Figure 4: Instruction tuning improves Mamba-2 ICL performance past Transformer LLMs. ALL LoRA models were instruction fine-tuned on the OpenHermes dataset for one epoch. Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks: HellaSwag, PIQA, Arc-E, Arc-C, and WinoGrande.

### v3

@@ 3.1.1 Divergence bounds for automatic mixed-precision and PEFT
Let $\epsilon_{t}^{u}=\max|\mathbf{u}_{t}-\mathbf{u}_{t}^{\prime}|$, i.e., the maximum, positive scalar-difference between the $t$th MPFT/PEFT and full fine-tuning input state. Similarly, let $\epsilon_{t}^{x}=\max|\bm{x}_{t}-\bm{x}_{t}^{\prime}|$ and $\epsilon^{*}=\max\bigcup_{t=0}^{T}\{\epsilon_{t}^{u},\epsilon_{t}^{x}\}$ (where we define $\mathbf{u}_{0}=0$). We thus have
$\max|F_{\theta}^{T}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{T}(\bm{x}_{0}^{\prime},\mathbf{u}_{1}^{\prime})|\in\mathcal{O}(\epsilon^{*}\exp{(T\zeta)}),$
for some scalar $\zeta\leq 0$. Thus, differences introduced by AMP and/or PEFT for Mamba models do not exponentially compound over discrete-time.
Theorem 4.
Let $\mathbf{y}_{t}^{\prime}$ be the trajectory of output states produced by $\bm{x}_{0}^{\prime},\dots,\bm{x}_{T}^{\prime}$ and $\mathbf{u}_{1}^{\prime},\dots,\mathbf{u}_{T}^{\prime}$ under MPFT and/or PEFT. Then we thus have
$\max|\mathbf{y}_{T}-\mathbf{y}_{T}^{\prime}|\in\mathcal{O}(\epsilon^{*}\exp{(T\zeta)}),$
for some scalar $\zeta\leq 0$. Thus, small changes to the output $\mathbf{y}_{t}$ resulting from MPFT and/or PEFT are exponentially non-increasing over discrete time.
The proof of Theorem 3 is available in Appendix C. Theorem 4 may be derived by directly applying the proof technique of Theorem 2 to the results of Theorem 3. Thus, the latent and output states of both Mamba and Mamba-2 models are stable to changes encountered during AMP and MFPT.
@@ 4 Experiments
4.1 Non-divergent MambaBlock dynamics
We explore the implications of Theorem 2 for arbitrary MambaBlocks. For $n$ random trials and various $\varepsilon$, we randomly initialize the inputs (i.e., $u_{t}$ for $t=1,\dots,T$) and parameters (i.e., $\mathbf{A},\mathbf{B}_{t},\mathbf{C}_{t},\bm{\Delta}_{t}$) of a MambaBlock with $T=2048$ and latent-variable dimension $d=64$.
For each trial and $\varepsilon$, we first calculate the MambaBlock outputs $\mathbf{y}_{1},\dots,\mathbf{y}_{T}$ (where $\mathbf{y}_{t}=F_{\theta}^{t}(\bm{x}_{0},\mathbf{u}_{1})$). We then perturb the initial states $(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)=(\bm{x}_{0}^{\prime},\mathbf{u}_{1}^{\prime})$ and calculate the subsequent perturbed model outputs $\mathbf{y}_{1}^{\prime},\dots,\mathbf{y}_{T}^{\prime}$ (where $\mathbf{y}_{t}^{\prime}=F_{\theta}^{t}(\bm{x}_{0}^{\prime},\mathbf{u}_{1}^{\prime})=F_{\theta}^{t}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}^{\prime}+\varepsilon)$). The maximum deviation is then calculated for each time point, i.e., $\max|y_{t}-y_{t}^{\prime}|=\max|F_{\theta}^{t}(\bm{x}_{0},\mathbf{u}_{1})-F_{\theta}^{t}(\bm{x}_{0}+\varepsilon,\mathbf{u}_{1}+\varepsilon)|$. Figure 1 shows both the mean maximum deviation per time averaged across trials, as well as the standard deviation, for $\varepsilon=0.1$.
As we can see, the maximum divergence starts out proportional to the change in initial conditions, but quickly decays and converges to the original output trajectory. Thus, changes to a MambaBlocks inputs do not compound over discrete time. Additional experiments for $\varepsilon\in\{0.01,0.001\}$ are available in Appendix D and demonstrate the same dynamics.
Figure 1: For 100 random MambaBlocks, initial latent and input states $(\bm{x}_{0},\mathbf{u}_{0})$ are deviated by $\varepsilon$. The resulting maximum deviation in output states is computed and shown to exponentially decrease over discrete time. $\mu$ denotes the maximum absolute deviation at discrete time $t$ averaged over all trials, while $\sigma$ denotes the respective standard deviation at each discrete time.
4.2 Non-divergent Mamba MPFT and PEFT
@@
We explore the implications of Theorem 4 for combinations of MPFT and PEFT on pretrained Mamba models. We focus on utilizing LoRA (Hu et al., 2021), which is arguably the most widely used PEFT framework for LLMs. Using the Alpaca dataset (Taori et al., 2023), Mamba 160M, 410M, and 790M models are fine-tuned for three epochs with a maximum sequence length of 512. We denote the targeting of all linear layers (ALL) for LoRA as ALL LoRA, the targeting of a subset of linear layers (SLL) for LoRA as SLL LoRA (i.e., $\bm{\Delta}_{t},\mathbf{B}_{t}$, and $\mathbf{C}_{t}$ parameters), and no adapters as Full (i.e., full fine-tuning).
@@
Each fine-tuning run occurred on a single Nvidia A10G GPU (24 GB total memory). To further limit extraneous numerical effects, the same batch size is used for all FP32, FP16, and BF16 experiments for a given model size. While this leads to hardware underutilization (i.e., non-saturated GPU memory for mixed-precision and LoRA experiments), this is necessary to guarantee no divergence is due to differences in parameter update schedules. For comparison, two Transformer-based LLM families of similar parameter counts are fine-tuned using the same experimental setup: Pythia (sizes 160M, 410M, and 1B) and OpenELM (Mehta et al., 2024) (sizes 270M and 450M). The training recipe for all models was adapted from Tunstall et al. (2023), with the AdamW_torch optimizer and a cosine annealing schedule. Further experimental details are available in Appendix F.
(a) Mean full-precision (FP32) divergence in MMLU performance.
(b) Mean full-precision (FP32) divergence in Winogrande performance.
Figure 2: Mean full-precision (FP32) divergence in performance for Mamba, Pythia, and OpenELM models. Models are fine-tuned over the Alpaca dataset using different combinations of MPFT and PEFT. Full fine-tuning (i.e., no PEFT adapters) is denoted as Full.
All models were evaluated using the LM evaluation harness from Eleuther AI (Gao et al., 2023). Model performance is measured as percent accuracy using the MMLU (Hendrycks et al., 2020) and Winogrande (Sakaguchi et al., 2021) datasets. The difference in model performance is reported as the mean divergence (i.e., absolute difference) between the original full-precision and respective mixed-precision model, averaged over {0, 1, 3, 5}-shot percent accuracy. Thus, a divergence greater than one denotes an average difference greater than one entire percentage of accuracy. We thus refer to a divergence greater than unity as a large deviation spike.
@@
Mean divergence results for all models are displayed in Figure 2(a) and Figure 2(b) for MMLU and Winogrande, respectively. Across mixed-precisions and adapter settings, Mamba displays smaller divergences than both Pythia and OpenELM models. E.g., for Winogrande evaluations, Alpaca fine-tuning with Mamba models is an average 2.9 and 2.4 times smaller in mean divergence than Pythia and OpenELM models, respectively. For MMLU evaluations, Alpaca fine-tuning with Mamba models is an average 2.6 times smaller in mean divergence than both Pythia and OpenELM models. Importantly, Mamba models do not exhibit large deviation spikes after fine-tuning, in contrast to both Pythia and OpenELM models.
We note that the results in Figure 2 display the fine-tuning of 72 LLMs and 576 few-shot evaluations. In Appendix H, these experiments are expanded to include the LIMA fine-tuning dataset (Zhou et al., 2024) as well as the standard deviations for all mean deviation results. Across all experiments, Mamba models are significantly more stable for MPFT/PEFT compared to Transformer-based LLMs. E.g., for MMLU evaluations, LIMA fine-tuning with Mamba models is an average 7 and 3.3 times smaller in mean divergence than Pythia and OpenELM models, respectively. For Winogrande evaluations, LIMA fine-tuning with Mamba models is an average 4.1 and 3.3 times smaller in mean divergence than Pythia and OpenELM models, respectively. Across the two fine-tuning datasets, two benchmarks, and six evaluated MPFT/PEFT combinations, Pythia and OpenELM produce 9 and 4 large deviation spikes, respectively, while Mamba produces zero.
4.3 Hyperparameter tuning robustness of Mamba models
We empirically demonstrate that Mamba models are robust to the choice of PEFT hyperparemters. Using the training recipe of  Tunstall et al. (2023) as a basis, we conduct an extensive hyperparameter search across the learning rate, LoRA dimension, and number of warmup steps. From the Cartesian-product of these three parameters, 150 hyperparameter configurations were sampled and used to fine-tune Mamba 370M over the Openhermes dataset (Teknium, 2024), which consists of 242,000 supervised samples. For comparison, Pythia 410M is similarly fine-tuned using the same set of 150 hyperparameter configurations.
(a) Mamba 370M
(b) Pythia 410M
Figure 3: Fine-tuning hyperparameter search for OpenHermes. Each point is a different hyperparameter configuration. SLL LoRA was used for both models. The $x$-axis is the learning rate, the $y$-axis is resulting MMLU 5-shot performance, bubble size is the LoRA dimension, and the color is the number of warmup steps $\in\{0,1\mbox{k},2\mbox{k}\}$.
The MMLU 5-shot performance for each of the 150 Mamba and Pythia fine-tuned models is displayed in Figure 3. Pythia 410M is capable of higher performance than Mamba 370M, where the average accuracy for the former and the latter are 26.5% and 24.8%, respectively. However, Mamba 370M is much more robust to the choice of hyperparameters, with a difference of 1.5% between the minimum (23.3%) and maximum (24.8%). In contrast, Pythia 410M fine-tuned models display a large performance difference of 4.7% between the minimum (22.9%) and maximum (27.6%).
@@ 4.3 Hyperparameter tuning robustness of Mamba models
4.4 Stable instruction tuning improves Mamba ICL for natural language tasks
@@
With both stable MPFT and PEFT, we determine the impact of instruction tuning Mamba and Mamba-2 models on ICL for natural language tasks. We instruction fine-tuned Mamba and Mamba-2 pretrained models using ALL LoRA, the OpenHermes dataset, and the training recipe described in Section 4 (with BF16). Zero and few-shot performance are evaluated using the same five standard natural language benchmarks used in Section 4.3. ICL performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot (AIPSS). For comparison, Pythia pretrained models are instruction fine-tuned using the same training recipe and ALL LoRA (i.e., all Pythia linear layers are adapted).
@@
Figure 4: Instruction tuning narrows the ICL gap between Mamba and Pythia, and creates a gap from Pythia to Mamba-2 models. ALL LoRA models were instruction tuned on the OpenHermes (Teknium, 2024) dataset for one epoch. Performance is reported as the average improvement percentage of {1, 3, 5}-shot versus 0-shot over five standard natural language benchmarks.
@@
However, after instruction tuning, all Mamba models larger than Mamba 130M consistently improve in ICL performance as the number of shots increase. Similarly, the majority of Mamba-2 models larger than Mamba-2 130M greatly improve in ICL performance. Thus, while pretrained Mamba and Mamba-2 models are only capable of 38% and 82% (respectively) of the AIPSS compared to similar pretrained Pythia models, instruction tuned Mamba and Mamba-2 models are capable of 81.5% and 133% of the AIPSS relative to similarly fine-tuned Pythia models.
In addition to Pythia, we compare instruction tuned Mamba models’ ICL abilities to other state-of-the Transformer LLMs of comparable sizes and pretraining token counts: OpenELM (Mehta et al., 2024) (sizes 270M, 450M, and 1.1B), TinyLlama 1.1B (Zhang et al., 2024), and OLMo 1.2B (Groeneveld et al., 2024). We additionally group models by parameter count, to understand ICL as an emergent behavior of fine-tuned Mamba models. To further explore the emergence of this ability for Mamba SSMs, we include recent Mamba-specific PEFT methods (Yoshimura et al., 2025). Displayed in Appendix G, pretrained SSMs and Transformers of parameter counts 270 million and less display slight or detrimental ICL abilities (i.e., few-shot performance is worse than zero-shot). For models of greater than 450 million parameters, the majority of SSMs and Transformers display positive ICL abilities, with the exception of Mamba 1.4B.